OpenStudy (anonymous):

An ice skating arena charges an admission fee for each child plus a rental fee for each pair of ice skates. John paid the admission fees for his six nephews and rented five pairs of ice skates. He was charged $32.00. Juanita paid the admission fees for her seven grandchildren and rented five pairs of ice skates. She was charged $35.25. What is the admission fee? What is the rental fee for a pair of skates?

OpenStudy (nurali):

C=price per child S=price per pair of ice skates John: 6C + 5S = 32.00 Juanita: 7C + 5S = 35.25 You can subtract one equation from the other to cancel out the S's and determine C: .... 6C + 5S = 32.00 .. -(7C + 5S = 35.25) _________________ .... 6C + 5S = 32.00 .. -7C - 5S = -35.25 _________________ -C + 0S = -3.25 ..........C=3.25 From this, it is determined that C=3.25. You can use this to determine S, by plugging it into one of the equations: 6C + 5S = 32.00 6(3.25) + 5S = 32.00 19.50 + 5S = 32.00 5S = 32.00 - 19.50 5S = 12.50 S = 12.50/5 S = 2.50 Now, it is determined that S = 2.50. You can verify this by plugging it into the other equation. 7C + 5S = 35.25 7(3.25) + 5S = 35.25 22.75 + 5S = 35.25 5S = 35.25 - 22.75 5S = 12.50 S = 12.50/5 S = 2.50 Again, S = 2.50. Therefore, it is correct.

OpenStudy (nurali):

Welcome to Openstudy.

OpenStudy (anonymous):

thank u

OpenStudy (nurali):

My Pleasure.